# Spatial-Aware Decision-Making with Ring Attractors in Reinforcement Learning Systems

### v2

@@ Abstract
This paper explores the integration of ring attractors, a mathematical model inspired by neural circuit dynamics, into the Reinforcement Learning (RL) action selection process. Serving as specialized brain-inspired structures that encode spatial information and uncertainty, ring attractors offer a biologically plausible mechanism to improve learning speed and accuracy in RL. They do so by explicitly encoding the action space, facilitating the organization of neural activity, and enabling the distribution of spatial representations across the neural network in the context of Deep Reinforcement Learning (DRL). For example, preserving the continuity between rotation angles in robotic control or adjacency between tactical moves in game-like environments. The application of ring attractors in the action selection process involves mapping actions to specific locations on the ring and decoding the selected action based on neural activity. We investigate the application of ring attractors by both building an exogenous model and integrating them as part of DRL agents. Our approach significantly improves state-of-the-art performance on the Atari 100k benchmark, achieving a 53% increase in performance across selected state-of-the-art baselines. Codebase available at RA-RL-8026.
Machine Learning, ICML
@@ 1 Introduction
This paper addresses the challenge of efficient action selection in Reinforcement Learning (RL), particularly in environments with spatial structures, from robotic manipulation where joint movements are coupled, to game-playing agents where tactical decisions might depend on positional awareness. We integrate ring attractors, a neural circuit model from neuroscience originally proposed by Zhang (1996) and subsequently discovered by Kim et al. (2017), into the RL framework. This approach provides a mechanism for uncertainty-aware decision-making in RL, thereby yielding more efficient and reliable learning in complex environments. Ring attractors offer a unique framework to represent spatial information in a continuous and stable manner (Sun et al., 2020). In a ring attractor network, neurons interconnect circularly, forming a loop with tuned connections (Blair et al., 2014). This configuration allows for robust and localized activation patterns, maintaining accurate spatial representations even with noise or perturbations. Applying ring attractors to the selection of RL actions involves mapping actions to specific ring locations and decoding the selected action based on neural activity. This spatial embedding proves advantageous for continuous action spaces, particularly in tasks such as robotic control and navigation (Rivero-Ortega et al., 2023). Ring attractors improve decision-making by exploiting spatial relations between actions, contributing to informed transitions between actions in sequential decision-making tasks in RL.
@@
  • We propose a novel approach for incorporating ring attractors into RL to incorporate spatial information and relations between actions. This spatial awareness significantly speeds up the learning rate of the RL agent. See Sections 3.1.2 and 4.1, respectively.
  • We utilize a Bayesian approach to build ring attractors input signals, encoding uncertainty estimation to drive the action selection process, showing further performance improvement compared to non-uncertainty based methods. See Sections 3.1.3 and 4.1, respectively.
  • We develop a reusable DL module based on recurrent neural networks that integrates ring attractors into DL-based RL agents. Allowing integration and comparison with existing Deep RL frameworks and baselines, enabling the adoption of our ring attractor approach across different RL models and tasks in applied domains. See 3.2 and 4.2, respectively.
2 Related Works
The integration of ring attractors into RL systems brings together neuroscience-inspired models and advanced machine learning techniques. Here, we review the literature on key areas that form the foundation of our RL research: spatial awareness in RL, biologically inspired RL approaches, and uncertainty quantification methods.
@@ 2.1 Spatial Awareness in RL
Incorporating spatial awareness into RL systems has improved performance on tasks with inherent spatial structure. Regarding relational RL, Zambaldi et al. (2019) introduced an approach using attention mechanisms to reason about spatial relations between entities in an environment. This method demonstrated improved sample efficiency and generalization in tasks that require spatial reasoning. On the topic of navigation, Mirowski et al. (2017) developed a DRL agent capable of navigating complex city environments using street-level imagery. Their approach incorporated auxiliary tasks, such as depth prediction and loop closure detection. Concerning explicit spatial representations, Gupta et al. (2017) proposed a cognitive mapping and planning approach for visual navigation, combining spatial memory with a differentiable neural planner. Similarly, Bapst et al. (2019) introduced a relational DRL framework using graph neural networks to capture spatial relations between objects. Although these approaches demonstrate the importance of spatial awareness in RL, they often lack the biological plausibility found in neural circuits.
@@ 2.2 Biologically Inspired Machine Intelligence
Biologically inspired approaches to RL seek to leverage insights from neuroscience to improve the efficiency, adaptability, and interpretability of RL algorithms. These methods often draw upon neural circuit dynamics and cognitive processes observed in biological systems. The work presented in (Banino et al., 2018) demonstrated that incorporating grid-like representations, inspired by mammalian grid cells, into RL agents improved performance in navigation tasks. Their work showed that these biologically inspired representations emerged naturally in agents trained on robotics tasks and transferred well to new environments. Similarly, Cueva & Wei (2018) showed that recurrent neural networks trained on navigation tasks naturally developed grid-like representations, suggesting a deep connection between biological and artificial navigation systems.  Singh et al. (2023) demonstrated how RL agents naturally develop insect-like behaviors and neural dynamics when solving complex spatial navigation tasks.  Wang et al. (2018) proposed a biologically inspired meta-RL algorithm that mimics the function of the prefrontal cortex and dopamine-based neuromodulation. Their approach demonstrated rapid learning and adaptation to new tasks, similar to the flexibility observed in biological learning systems.
@@ 2.3 Uncertainty Quantification
Regarding exploration strategies, Osband et al. (2016) introduced bootstrapped Deep Q-networks (DQNs), addressing exploration by leveraging uncertainty in Q-value estimates by training multiple DQNs with shared parameters. Building on this theme, Burda et al. (2018) proposed random network distillation (RND), measuring uncertainty by comparing predictions between a target network and a randomly initialized network. For efficient uncertainty quantification, Durasov et al. (2020) and Bykovets et al. (2022) presented a novel ‘masksemble’ approach, applying masks across the input batch during the forward pass to generate diverse predictions. Addressing risk assessment in non-stationary environments, Jain et al. (2021) described a method to analyze sources of lack of knowledge by adding a second Bayesian model to predict algorithmic action risks, particularly relevant for multi-agent RL (MARL) systems.  Kutschireiter et al. (2023) developed a Bayesian ring attractor that outperforms conventional ring attractors by dynamically adjusting its activity based on evidence quality and uncertainty. In the context of individual treatment effects, Lee et al. (2020) performed uncertainty quantification (UQ) using an exogenously prescribed algorithm, making the method agnostic to the underlying recommender algorithm. Azizzadenesheli et al. (2018) developed a Bayesian approach for RL in episodic high-dimensional Markov decision processes (MDPs). They introduced two novel algorithms: LINUCB and LINPSRL. These algorithms achieve significant improvements in sample efficiency and performance by incorporating uncertainty estimation into the learning process. In the context of DRL, called Bayesian Deep Q-networks, BDQNs (Azizzadenesheli et al., 2018), incorporates efficient Thompson sampling and Bayesian linear regression at the output layer to factor uncertainty estimation in the action-value estimates.
In summary, the literature reveals a growing interest in incorporating spatial awareness, biological inspiration, and UQ into RL systems. However, there remains a gap in integrating these elements into a cohesive framework. Our work on ring attractors aims to bridge this gap by providing a biologically plausible model that inherently captures spatial relations and can be extended to handle uncertainty, potentially leading to more robust and efficient RL agents.
@@ 3 Methodology
In this section, we describe two main methods: an exogenous ring attractor model using continuous-time recurrent neural networks (CTRNNs) and a DL-based ring attractor integrated into the RL agent. While the CTRNN model validates the theoretical principles, the DL architecture eases deployment and adoption in existing DRL frameworks. Both leverage the ring attractors’ spatial encoding capabilities to enhance action selection and performance. In the CTRNN model, we detail the ring attractor architecture, Section 3.1.1; dynamics and implementation, Section 3.1.2; and uncertainty injection, Section 3.1.3. CTRNNs are used for their ability to model continuous neural dynamics and maintain stable attractor states (Beer, 1995). The integrated DL approach offers end-to-end training for efficiency and scalability, detailed in Section 3.2. Ring attractors in RL maintain stable spatial information representations, preserving action relations lost in traditional flattened action spaces. This circular spatial representation potentially yields smoother policy gradients and more efficient learning in spatial tasks, attributed to the ring attractors’ ability to maintain a stable representation of spatial information.
3.1 Continuous–Time RNN Ring Attractor Model
@@ 3.1.1 Ring Attractor Architecture
Ring attractors commonly consist of a configuration of excitatory and inhibitory neurons arranged in a circular pattern. We can model the dynamics of the ring using the Touretzky ring attractor network (Touretzky, 2005). In this model, each excitatory neuron establishes connections with all other excitatory neurons, and an inhibitory neuron is placed in the middle of the ring with equal weighted connections to all excitatory neurons.
@@
$x_{n}(K_{i})=\sum_{i=1}^{I}x^{i}_{n}(\alpha_{n})=\sum_{i=1}^{I}{\frac{K_{i}}{\sqrt{2\pi\sigma_{i}}}\exp\left(-\frac{(\alpha_{n}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right)}$ (1)
@@
$\frac{dv_{n}}{dt}\approx\frac{\Delta v_{n}}{\Delta t}=\frac{v_{n+\Delta t}-v_{n}}{\Delta t}=\frac{f(x_{n}+\epsilon_{n}+\eta_{n})}{\tau}-v_{n}$ (2)
@@
$\frac{du}{dt}\approx\frac{\Delta u}{\Delta t}=\frac{u_{+\Delta t}-u}{\Delta t}=\frac{f(\epsilon_{n}+\eta_{n})}{\tau}-u$ (3)
@@ 3.1.2 Ring Attractor as Behavior Policy in RL
Excitatory neurons’ action input signal: We begin by reformulating the input function for a target excitatory neuron $n$ from Eq. (1). The key modification is setting the scale factor $K_{i}$ to the Q-value $Q(s,a)$ of the state-action pair $(s,a)$, that is $K_{i}=Q(s,a)$.
Figure 1: Ring attractor Touretzky representation: Circular arrangement of excitatory neurons (N0-N7) with excitatory connections and central inhibitory neuron. Four input signals shown as colored gradients. Overall activation depicted by red outline. Includes connection weights and input signal parameters, illustrating ring attractor dynamics.
This formulation is represented in Fig. 1 and described in Eq. (5) ensures that actions with higher estimated values are given more weight in the ring attractor dynamics, naturally biasing the network towards more valuable actions. The orientation of the signal within the ring attractor is determined by the direction of movement in the action space. We represent this as $\mu_{i}=\alpha_{a}(a)$, where $\alpha_{a}(a)$ is the angle corresponding to the action $a$ in the circular action space. We define our circular action space $\mathcal{A}$ as a subset of $\mathbb{R}^{2}$, where each action $a\in\mathcal{A}$ is represented by a point on the unit circle. The function $\alpha:\mathcal{A}\rightarrow[0,2\pi)$ maps each action to its corresponding angle on this circle, and $\alpha_{n}$ which presents the preference for the orientation of the neuron $n$ in space. To account for uncertainty in our value estimates, we incorporate the variance of the estimated value for each action into our model: $\sigma_{i}=\sigma_{a}$.
This allows the network to represent not just the expected value of actions, but also our confidence in those estimates. Combining these elements, we arrive at the following equation for the action signal $x(Q)$:
$x_{n}(Q)=\sum_{a=1}^{A}{\frac{Q(s,a)}{\sqrt{2\pi\sigma_{a}}}\exp\left(-\frac{1}{2}\frac{(\alpha_{n}-\alpha_{a}(a))^{2}}{\sigma_{a}^{2}}\right)}$ (5)
@@
$\frac{dv_{n}}{dt}=\frac{1}{\tau}$ $\left(max\left(0,\left(\sum_{m=1}^{m=N}w^{(E\xrightarrow{}E)}_{m,n}v_{m}+\right.\right.\right.$
$\left.\left.\left.x_{n}(Q)+w^{(I\xrightarrow{}E_{u})}u\right)\right)\right)-v_{n}$ (6)
@@
$\frac{du}{dt}=\frac{1}{\tau}\left(max\left(0,\left(u+\sum_{n=1}^{N}w^{(E_{n}\rightarrow I)}_{n}v_{n}\right)\right)\right)-u,$ (7)
where the self-inhibition term $u$ and excitatory input $v_{n}$ from each excitatory neuron $n$ from Eq. (3.1.2) updates the activation of the inhibitory neuron $\frac{du}{dt}$. Eq. (3.1.2) and Eq. (7) collectively describe the dynamics of the ring attractor network, capturing the interplay between excitatory and inhibitory neurons, external inputs, and synaptic connections.
To translate the ring attractor’s output into an action in the 2D space, we use the following equation:
$\text{action}=argmax\{\mathbf{V}\}\frac{N^{(A)}}{N^{(E)}},$ (8)
where $n\in\{1,...,N^{(E)}\}$, $N^{(E)}$ is the number of excitatory neurons in the ring attractor, $N^{(A)}$ is the number of discrete actions in the action space $\mathcal{A}$, and the activation of excitatory neurons around the ring $\mathbf{V}=[v_{1},v_{2},...,v_{N^{(E)}}]$.
This equation assumes that both the neurons in the ring attractor and the actions in the action space are uniformly distributed. This approach allows for nuanced action selection that takes into account both the spatial relations between actions and their estimated values. A visualization of the ring is presented in Fig. 1.

### v3

@@ Abstract
Ring attractors, mathematical models inspired by neural circuit dynamics, provide a biologically plausible mechanism to improve learning speed and accuracy in Reinforcement Learning (RL). Serving as specialized brain-inspired structures that encode spatial information and uncertainty, ring attractors explicitly encode the action space, facilitate the organization of neural activity, and enable the distribution of spatial representations across the neural network in the context of Deep Reinforcement Learning (DRL). These structures also provide temporal filtering that stabilizes action selection during exploration, for example, by preserving the continuity between rotation angles in robotic control or adjacency between tactical moves in game-like environments. The application of ring attractors in the action selection process involves mapping actions to specific locations on the ring and decoding the selected action based on neural activity. We investigate the application of ring attractors by both building an exogenous model and integrating them as part of DRL agents. Our approach significantly improves state-of-the-art performance on the Atari 100k benchmark, achieving a 53% increase in performance over selected baselines.
@@ 1 Introduction
This paper addresses the challenge of efficient action selection in Reinforcement Learning (RL), particularly in environments with spatial structures, ranging from robotic manipulation where joint movements are coupled, to game-playing agents where tactical decisions might depend on positional awareness. We integrate ring attractors, a neural circuit model from neuroscience originally proposed by AttractorsFirstIntuition and later experimentally validated by RA_Science , into the RL framework. This approach provides a mechanism for uncertainty-aware decision-making in RL, thereby yielding more efficient and reliable learning in complex environments. Ring attractors offer a unique framework to represent spatial information in a continuous and stable manner RA_Evidence . In a ring attractor network, neurons interconnect circularly, forming a loop with tuned connections RA_Biological_Plausability . This configuration allows for robust and localized activation patterns, maintaining accurate spatial representations even in the presence of noise or perturbations. Applying ring attractors to the selection of RL actions involves mapping actions to specific ring locations and decoding the selected action based on neural activity. This spatial embedding proves advantageous for continuous action spaces, particularly in tasks such as robotic control and navigation RA_Robot_Navigation . Ring attractors improve decision-making by exploiting spatial relations between actions, contributing to informed transitions between actions in sequential decision-making tasks in RL. While many action spaces possess inherent topological structure todorov2012mujoco , traditional neural networks represent actions as orthogonal vectors that ignore these relationships. Ring attractors bridge this gap by providing a neural substrate that preserves spatial ordering and similarity between actions. This is particularly relevant given recent advances showing that spatial structure in representations improves sample efficiency spatial_navigation ; Relational_bias and generalization Structured_representations ; RL_explicit_representations , yet these benefits have not been systematically applied to action selection mechanisms themselves.
@@
  • We propose a novel approach for integrating ring attractors into RL, incorporating spatial information, temporal filtering, and relations between actions. This spatial awareness significantly speeds up the learning rate of the RL agent. See Sections 3.1.2 and 4.1.
  • We utilize a Bayesian approach to build ring attractors input signals, encoding uncertainty estimation to drive the action selection process, showing further performance improvement. See Sections 3.1.3 and 4.1.
  • We develop a reusable DL module based on recurrent neural networks that incorporates ring attractors into DRL agents. This allows integration and comparison with existing DRL frameworks and baselines, enabling the adoption of our ring attractor approach across different RL models and tasks in applied domains. See Sections 3.2 and 4.2.
Codebase available at https://github.com/marcosaura/RA_RL.
2 Related Research
The integration of ring attractors into RL systems brings together neuroscience-inspired models and advanced machine learning techniques. In this section, we review the literature on key areas that form the foundation of our RL research: spatial awareness in RL, biologically inspired RL approaches, and uncertainty quantification methods.
@@ 2.1 Spatial Awareness in RL
Incorporating spatial awareness into RL systems has improved performance on tasks with inherent spatial structure. Regarding relational RL, Relational_bias introduced an approach using attention mechanisms to reason about spatial relations between entities in an environment. This method demonstrated improved sample efficiency and generalization in tasks that require spatial reasoning. On the topic of navigation, spatial_navigation developed a DRL agent capable of navigating complex city environments using street-level imagery. Their approach incorporated auxiliary tasks, such as depth prediction and loop closure detection. Concerning explicit spatial representations, RL_explicit_representations proposed a cognitive mapping and planning approach for visual navigation, combining spatial memory with a differentiable neural planner. Similarly, Structured_representations introduced a relational DRL framework using graph neural networks to capture spatial relations between objects. Although these approaches demonstrate the importance of spatial awareness in RL, they rely on emergent learning to develop spatial understanding through comprehensive architectural additions, which require extensive training to discover action relationships. In contrast, our ring attractor approach explicitly encodes spatial structure in the action space, providing spatial inductive bias that accelerates learning rather than relying on the RL agent to autonomously discover all critical action space relationships, which may result in suboptimal or incomplete spatial understanding.
@@ 2.2 Biologically Inspired Machine Intelligence
Biologically inspired approaches to RL seek to leverage insights from neuroscience to improve the efficiency, adaptability, and interpretability of RL algorithms. These methods often draw upon neural circuit dynamics and cognitive processes observed in biological systems. A notable example is the work in wang2002probabilistic , which demonstrated how neural attractor states implement probabilistic choices through recurrent network dynamics. The work presented in navigation_grid demonstrated that incorporating grid-like representations, inspired by mammalian grid cells, into RL agents improved performance in navigation tasks. Their work showed that these biologically inspired representations emerged naturally in agents trained on robotics tasks and transferred well to new environments. Similarly, emerging_grid_representations showed that recurrent neural networks trained on navigation tasks naturally developed grid-like representations, suggesting a deep connection between biological and artificial navigation systems. In RL, grid cells and ring attractors serve distinct computational roles. Grid cells encode spatial state information (the agent’s position), through hexagonal firing patterns, functioning as fixed basis functions for value functions and spatial generalization navigation_grid ; emerging_grid_representations . Ring attractors encode persistent states, directional and spatial bias, through recurrent dynamics that actively maintain information over time khona2022attractor . While grid cells provide emergent spatial input features, ring attractors are dynamical components that provide intrinsic spatial encoding for memory states kilpatrick2013wandering , that here we translate to encoding of the action space. PlumesNav demonstrated how RL agents naturally develop insect-like behaviors and neural dynamics when solving complex spatial navigation tasks. rapid_learnig proposed a biologically inspired meta-RL algorithm that mimics the function of the prefrontal cortex and dopamine-based neuromodulation. Their approach demonstrated rapid learning and adaptation to new tasks, similar to the flexibility observed in biological learning systems. However, these approaches primarily focus on mimicking neural representations or learning dynamics, whereas ring attractors specifically encode spatial relationships in the action space itself, providing both biological plausibility and direct spatial awareness for action selection.
@@ 2.3 Uncertainty Quantification
Regarding exploration strategies, BootstrappedDQN introduced bootstrapped Deep Q-networks (DQNs), addressing exploration by leveraging uncertainty in Q-value estimates by training multiple DQNs with shared parameters. Building on this theme, RandomNetDistillation proposed random network distillation (RND), measuring uncertainty by comparing predictions between a target network and a randomly initialized network. For efficient uncertainty quantification, Masksemble and Masksemble_Support_Article presented a novel “masksemble” approach, applying masks across the input batch during the forward pass to generate diverse predictions. Addressing risk assessment in non-stationary environments, Epistemic_Uncertainty_Prediction described a method to analyze sources of lack of knowledge by adding a second Bayesian model to predict algorithmic action risks, particularly relevant for multi-agent RL (MARL) systems. BRA developed a Bayesian ring attractor that outperforms conventional ring attractors by dynamically adjusting its activity based on evidence quality and uncertainty. In the context of individual treatment effects, Grouping_Uncertainty_Sources performed uncertainty quantification (UQ) using an exogenously prescribed algorithm, making the method agnostic to the underlying recommender algorithm. BDQN developed a Bayesian approach for RL in episodic high-dimensional Markov decision processes (MDPs). They introduced two novel algorithms: LINUCB and LINPSRL. These algorithms achieve significant improvements in sample efficiency and performance by incorporating uncertainty estimation into the learning process. In the context of DRL, Bayesian Deep Q-networks (BDQNs) BDQN incorporate efficient Thompson sampling and Bayesian linear regression at the output layer to factor uncertainty estimation in the action-value estimates. While these methods treat uncertainty estimation as a separate module applied to existing architectures, our ring attractor approach integrates uncertainty through the variance parameters of Gaussian input signals, making uncertainty awareness an intrinsic part of the spatial action representation.
In summary, the literature reveals a growing interest in incorporating spatial awareness, biological inspiration, and UQ into RL systems. However, there remains a gap in integrating these elements into a cohesive framework that simultaneously provides biological plausibility, explicit spatial encoding, and natural uncertainty handling within a single neural architecture. Our work on ring attractors aims to bridge this gap by providing a biologically plausible model that inherently captures spatial relations and can be extended to handle uncertainty, potentially leading to more robust and efficient RL agents.
@@ 3 Methodology
Ring attractors represent a computational principle where neurons are arranged in a circular topology to maintain spatial representations. Originally proposed by AttractorsFirstIntuition for encoding heading direction and empirically discovered by RA_Science in Drosophila, these networks exhibit circular organization where head direction cells encode different spatial directions (RA_evidence_mammals, ). For our methodology, ring attractors integrate multiple cues through distance-weighted connections (RA_Sensor_Fusion_Evidence, ; Cue_Control_RA, ) and maintain representational stability through excitatory-inhibitory dynamics (RA_evidence_mammals_second, ). This bioinspired structure has been repurposed in this line of research to explicitly encode the action space during the decision-making process. In this section, we describe two main methods: an exogenous ring attractor model using continuous-time recurrent neural networks (CTRNNs) and a DL-based ring attractor integrated into the RL agent. While the CTRNN model validates the theoretical principles, the DL architecture eases deployment and adoption in existing DRL frameworks. Both leverage the ring attractors’ spatial encoding capabilities to enhance action selection and performance. In the CTRNN model, we detail the ring attractor architecture, Section 3.1.1; dynamics and implementation, Section 3.1.2; and uncertainty injection, Section 3.1.3. CTRNNs are used for their ability to model continuous neural dynamics and maintain stable attractor states (CTRN, ). The integrated DL approach offers end-to-end training for efficiency and scalability, detailed in Section 3.2. Ring attractors in RL maintain stable spatial information representations, preserving action relations lost in traditional flattened action spaces. This circular spatial representation potentially yields smoother policy gradients and more efficient learning in spatial tasks, attributed to the ring attractors’ ability to maintain a stable representation of spatial information.
3.1 Continuous-Time RNN Ring Attractor Model
@@ 3.1.1 Ring Attractor Architecture
Ring attractors commonly consist of a configuration of excitatory and inhibitory neurons arranged in a circular pattern. We can model the dynamics of the ring using the Touretzky ring attractor network (Touretzky_ring, ). In this model, each excitatory neuron establishes connections with all other excitatory neurons, and an inhibitory neuron is placed in the middle of the ring with equal weighted connections to all excitatory neurons.
@@
$x_{n}(\alpha_{n})=\sum_{i=1}^{I}x^{i}_{n}(\alpha_{n})=\sum_{i=1}^{I}{\frac{K_{i}}{\sqrt{2\pi\sigma_{i}}}\exp\left(-\frac{(\alpha_{n}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right)}$ (1)
@@
$\frac{\mathrm{d}v_{n}}{\mathrm{d}t}\approx\frac{\Delta v_{n}}{\Delta t}=\frac{v_{n+\Delta t}-v_{n}}{\Delta t}=\frac{f(x_{n}+\epsilon_{n}+\eta_{n})}{\tau}-v_{n}$ (2)
@@
$\frac{\mathrm{d}u}{\mathrm{d}t}\approx\frac{\Delta u}{\Delta t}=\frac{u_{+\Delta t}-u}{\Delta t}=\frac{f(\epsilon_{n}+\eta_{n})}{\tau}-u$ (3)
@@ 3.1.2 Ring Attractor as Behavior Policy in RL
Excitatory neurons’ action input signal: We begin by reformulating the input function for a target excitatory neuron $n$ from Eq. (1). To integrate the ring attractor with RL, we reformulate the input signals by mapping $K_{i}$ to the Q-value $Q(s,a)$, effectively substituting input signal index $i\in I$ with actions $a\in A$. In other words, the key modification is setting the scale factor $K_{i}$ to the Q-value $Q(s,a)$ of the state-action pair $(s,a)$, that is $K_{i}=Q(s,a)$.
This formulation is represented in Fig. 1 and described in Eq. (5) ensures that actions with higher estimated values are given more weight in the ring attractor dynamics, naturally biasing the network towards more valuable actions. The orientation of the signal within the ring attractor is determined by the direction of movement in the action space. We represent this as $\mu_{i}=\alpha_{a}(a)$, where $\alpha_{a}(a)$ is the angle corresponding to the action $a$ in the circular action space. We define our circular action space $\mathcal{A}$ as a subset of $\mathbb{R}^{2}$, where each action $a\in\mathcal{A}$ is represented by a point on the unit circle. The function $\alpha:\mathcal{A}\rightarrow[0,2\pi)$ maps each action to its corresponding angle on this circle, and $\alpha_{n}$ which presents the preference for the orientation of the neuron $n$ in space. To account for uncertainty in our value estimates, we incorporate the variance of the estimated value for each action $\sigma_{i}=\sigma_{a}$.
This allows the network to represent not just the expected value of actions, but also the confidence in those estimates. Combining these, we arrive at the following equation for the action signal $x_{n}(Q)$:
$x_{n}(Q(s,a))=\sum_{a=1}^{A}{\frac{Q(s,a)}{\sqrt{2\pi\sigma_{a}}}\exp\left(-\frac{(\alpha_{n}-\alpha_{a}(a))^{2}}{2\sigma_{a}^{2}}\right)}$ (5)
@@
$\frac{\mathrm{d}v_{n}}{\mathrm{d}t}=\frac{1}{\tau}\left(max\left(0,\left(\sum_{m=1}^{m=N}w^{(E\xrightarrow{}E)}_{m,n}v_{m}+x_{n}(Q)+w^{(I\xrightarrow{}E_{u})}u\right)\right)\right)-v_{n}$ (6)
@@
$\frac{\mathrm{d}u}{\mathrm{d}t}=\frac{1}{\tau}\left(max\left(0,\left(u+\sum_{n=1}^{N}w^{(E_{n}\rightarrow I)}_{n}v_{n}\right)\right)\right)-u,$ (7)
Figure 1: Ring attractor Touretzky representation: Circular arrangement of excitatory neurons (N0-N7) and central inhibitory neuron. Four input signals shown as colored gradients. Overall activation depicted by red outline. Includes connection weights and input signal parameters, illustrating the final ring attractor dynamics state from Eq. (8).
where the self-inhibition term $u$ and excitatory input $v_{n}$ from each excitatory neuron $n$ from Eq. (6) updates the activation of the inhibitory neuron $\frac{\mathrm{d}u}{dt}$. Eq. (6) and Eq. (7) collectively describe the dynamics of the ring attractor network, capturing the interplay between excitatory and inhibitory neurons, external inputs, and synaptic connections. To translate the ring attractor’s output into an action in the 2D space, we use the following equation:
$\text{action}=\mathrm{argmax}\{\mathbf{V}\}\frac{N^{(A)}}{N^{(E)}},$ (8)
where $n\in\{1,...,N^{(E)}\}$, $N^{(E)}$ is the number of excitatory neurons in the ring attractor, $N^{(A)}$ is the number of discrete actions in the action space $\mathcal{A}$, and the activation of excitatory neurons around the ring $\mathbf{V}=[v_{1},v_{2},...,v_{N^{(E)}}]$. This equation assumes that both the neurons in the ring attractor and the actions in the action space are uniformly distributed. This approach allows for nuanced action selection that takes into account both the spatial relations between actions and their estimated values. A visualization of the ring is presented in Fig. 1.